OpenStudy (anonymous):

please integrate sine^3x ???

OpenStudy (anonymous):

\[\int\limits_{}^{}\sin (3x) \ dx\] Perform a substitution, where \[u = 3x \ \ du = 3 \ dx \ \ \frac{1}{3}du = 1 \ dx\] Then you have an integral \[\frac{1}{3} \int\limits\limits_{}^{}\sin (u) \ du = -\frac{1}{3}\cos(u) + c\] Substitute u = 3x back in and you'll have your answer \[-\frac{1}{3}\cos(3x) + c\]